# On the complex zeros and the computational complexity of approximating the reliability polynomial

### v1

@@ 1.1. Our contributions
(1.2) $S(G;p)=\sum_{\begin{subarray}{c}A\subseteq E\\ (V,A)\text{ $s-t$ split}\end{subarray}}(1-p)^{|A|}p^{|E|-|A|}.$
Alternatively, let $\hat{G}$ be the graph obtained from $G$ by merging the vertices $s$ and $t$ (and any edge between $s$ and $t$ becoming a loop). Then $S(G;p)=R(\hat{G};p)-R(G;p)$. Next we define the effective edge interaction of $G$ at $p$ as
(1.3) $y_{G}(p):=(1-p)\frac{S(G;p)}{R(G;p)}+1$
@@
and refer to these sets as the zero-locus, density locus and activity locus of $\mathcal{H}_{G_{0}}$ respectively. We moreover define the real analogues of these loci, restricting $p$ to be real, and denote these by $\mathcal{Z}_{G_{0}}^{\mathbb{R}},\mathcal{D}_{G_{0}}^{\mathbb{R}}$ and $\mathcal{A}_{G_{0}}^{\mathbb{R}}$ respectively, where for the real activity locus we also require that $\hat{y}_{G}(p)<-1$ and of course disregard the requirement that $\hat{y}_{G}(p)\notin\mathbb{R}$.
Our first main result relate the different loci defined above and will be proved in Section 3.
@@
$\overline{\mathcal{Z}_{G_{0}}}=\overline{\mathcal{A}_{G_{0}}}\quad\text{and}\quad\mathcal{A}_{G_{0}}\setminus\{p\mid R(G_{0};p)=0\}\subseteq\mathcal{D}_{G_{0}}\subseteq\mathcal{A}_{G_{0}}.$
@@
Name: $p$-Abs-Approx-Planar-Reliability
@@
Name: $p$-Arg-Approx-Planar-Reliability
@@
We define the problems $p$-Abs-Approx-Reliability and $p$-Arg-Approx-Reliability in a similar way, putting no restrictions on the input graph $H$. Note that if $p$ is a real number finding a $0.25$-arg approximation of $R(H;p)$ is equivalent to finding its sign (assuming $R(H;p)\neq 0$).
@@
Let $G_{0}$ two-terminal. Then for any algebraic number $p\in\mathcal{D}_{G_{0}}\cup\mathcal{D}^{\mathbb{R}}_{G_{0}}$ such that $R(G_{0};p)S(G_{0};p)\neq 0$,
  • the problems $p$-Abs-Approx-Reliability and $p$-Arg-Approx-Reliability are #P-hard;
  • if additionally $G_{0}$ is planar with its two terminals on the same face, then the problems $p$-Abs-Approx-Planar-Reliability and $p$-Arg-Approx-Planar-Reliability are #P-hard.
@@
For each algebraic number $p\in\mathbb{D}\setminus[0,1)$ both the problems $p$-Abs-Planar-Reliability and $p$-Arg-Planar-Reliability are #P-hard.
@@ 2.1. Graph operations and reliability
$R(G_{1}\parallel G_{2};p)$ $=R(G_{1};p)S(G_{2};p)\!+\!\!R(G_{2};p)S(G_{1};p)\!+\!\!R(G_{1};p)R(G_{2};p),$
$S(G_{1}\bowtie G_{2};p)$ $=R(G_{1};p)S(G_{2};p)+R(G_{2};p)S(G_{1};p).$
@@
Let $p\in[0,1]$; by definition, this makes $R(G;p)$ and $S(G;p)$ the probabilities that $G$ remains connected/becomes an sts-t split respectively. We omit the variable $p$ for ease of notation. Then we have the following.
@@ 2.2. Edge interactions
$y_{G_{1}\bowtie G_{2}}$ $=\frac{(1-p)S(G_{1}\bowtie G_{2})}{R(G_{1}\bowtie G_{2})}+1$
$=\frac{(1-p)\left(R(G_{1})S(G_{2})+R(G_{2})S(G_{1})\right)}{R(G_{1})R(G_{2})}+1$
$=\frac{(1-p)R(G_{2})S(G_{1})}{R(G_{1})R(G_{2})}+\frac{(1-p)R(G_{1})S(G_{2})}{R(G_{1})R(G_{2})}+1$
$=\left(\frac{(1-p)S(G_{1})}{R(G_{1})}+1\right)+\left(\frac{(1-p)S(G_{2})}{R(G_{2})}+1\right)-1$
@@
$\hat{y}_{G_{1}\parallel G_{2}}$ $=\frac{R(G_{1}\parallel G_{2})}{S(G_{1}\parallel G_{2})}+1$
$=\frac{S(G_{1}\parallel G_{2})+R(G_{1}\parallel G_{2})}{S(G_{1}\parallel G_{2})}$
$=\frac{S(G_{1})S(G_{2})+R(G_{1})S(G_{2})+R(G_{2})S(G_{1})+R(G_{1})R(G_{2})}{S(G_{1})S(G_{2})}$
$=\frac{(S(G_{1})+R(G_{1}))(S(G_{2})+R(G_{2}))}{S(G_{1})S(G_{2})}$
@@ 2.3. Gadgets
Take a two-terminal graph $(G,s,t)$, which we refer to as a gadget, and another graph $H$. We can substitute the gadget into an edge $e$ of $H$ by removing $e$, adding a copy of $G$, and identifying the source and sink of $G$ with the endpoints of $e$; we denote this operation as $H(G)_{e}$. See Figure 2 below for an example. While technically it matters which endpoint of $e$ is identified with $s$, for the purpose of the reliability polynomial it does not matter as we show below and therefore we don’t specify the choice of identification.
@@
$S(H(G)_{e};p)=S(G;p)S(H\setminus e;p)+R(G;p)S(H/e;p).$
@@
Let $H=(V(H),E(H))$ be a graph and $G=(V(G),E(G))$ a gadget; let $H(G)_{e}=(V,E)$. Consider the structure of a connected edge subgraph of $H(G)_{e}$; every vertex of $G$ has to be connected to at least one of $s,t$ to be connected to the vertices of $H$. A subset $A\subset E$ naturally decomposes into $E_{H}\cup E_{G}$, with $A_{H}\subseteq E(H)$, $A_{G}\subseteq E(G)$ and $E_{H}\cap E_{G}=\emptyset$. Then a spanning subgraph $(V,A)$ of $H(G)_{e}$ is connected if and only if one of the following happens:
@@
$\frac{1-p}{R(G;p)}R(H(G)_{e};p)$ $=\frac{1-p}{R(G;p)}\left(S(G;p)R(H\setminus e;p)+R(G;p)R(H/e;p)\right)$
$=(1-p)\frac{S(G;p)}{R(G;p)}R(H\setminus e;p)+R(H;p)-pR(H\setminus e;p)$
@@
(2.3) $R(H(G);p)$ $=\left(R(G;p)+S(G;p)\right)^{|E(H)|}R\left(H;\frac{1}{\hat{y}_{G}(p)}\right)\ \ \text{and}$
$S(H(G);p)$ $=\left(R(G;p)+S(G;p)\right)^{|E(H)|}S\left(H;\frac{1}{\hat{y}_{G}(p)}\right),$
from which the statement of the lemma directly follows. Observe that $R(G;p)$ is zero as a polynomialproof if and only if $G$ is disconnected. We can then assume that $p\in[0,1]$ is such that $R(G;p)\neq 0$. We work by induction on the number of edges of $H$, starting with the case $|E(H)|=1$. We may assume $H=K_{2}$ (for otherwise the statement is trivial). Since $K_{2}(G)=G$, it is not difficult to see that the desired statement holds. Let us next assume that $|E(H)|>1.$ Let us view $H(G)$ as $H^{\prime}(G)_{e}$ for some edge $e$ of $H$. (Here $H^{\prime}$ is the graph obtained from $H(G)$ be replacing the copy of $G$ on $e$ by $K_{2}$. by Lemma 2.5 we then have
$R(H(G)$ $;p)=\,R(G;p)R((H/e)(G);p)+S(G;p)R((H\setminus e)(G);p)$
$=$ $\,\left(R(G;p)+S(G;p)\right)\!\!\Bigg(\!\!\!\left(1-\frac{1}{\hat{y}_{G}(p)}\right)R((H/e)(G);p)+\frac{1}{\hat{y}_{G}(p)}R((H\setminus e)(G);p)\!\!\Bigg).$
Since $H/e,H\setminus e$ both have fewer edges than $H$, the statement follows by induction and the deletion contraction recurrence. The same steps work if we substitute $R(H(G);p)$ with $S(H(G);p)$. ∎
@@ 2.4. Multivariate reliability and split reliability
In this section we will define the multivariate analogue of the reliability and split reliability of a graph. As we will see this will help us to derive formula for the the general edge replacement, where we replace each edge $e_{i}$ of $H$ with different two-terminal graphs $G_{i}$. For a graph $G=(V,E)$ with edges $\{e_{1},\dots,e_{m}\}$ let $p_{1},\dots,p_{m}$ variables. Then the multivariate reliability polynomial is defined as
@@
Let $H$ be a two-terminal graph with edges $\{e_{1},\dots,e_{m}\}$ and let $G_{1},\dots,G_{m}$ be two-terminal graphs. Let $H(G_{1},\dots,G_{m})$ be a graph obtained from $H$ such that we replace the edge $e_{i}$ with the two-terminal graph $G_{i}$ (i.e. we delete the edge $e_{i}=(u,v)$ and we identify the source (resp. target) of $G_{i}$ with $u$ (resp. $v$). Then
@@ 2.5. Normal families and the Montel-Carathèodory Theorem
A family $\mathcal{F}$ of meromorphic functions on an open set $U\subseteq\hat{\mathbb{C}}$ is called normal if every sequence $\{f_{n}\}\subseteq\mathcal{F}$ contains a subsequence which converges spherically uniformly on compact subsets of $U$.
@@
Let $(f_{n})$ be a sequence of meromorphic functions on $U$ which converges spherically uniformly on compact subsets to $f$. Then $f$ is either a meromorphic function on $U$ or identically equal to $\infty$.
@@ 3.1. Relating the activity-locus and the zero-locus
We begin by stating a lemma on the values of the virtual edge interaction (which we recall are defined in (1.4)) at reliability zeros.
@@
Let $p\in\mathbb{C}$ and let $G_{0}$ be a two-terminal graph such that $S(G_{0};p)\neq-R(G_{0};p)$. Then the following are equivalent:
@@
    $\hat{y}_{G}(p)\in\{\omega,\omega^{2},1\}$, $S(G;p)\neq 0$ for some $G\in\mathcal{H}_{G_{0}}$ and $\omega$ a primitive third root of unity.
@@
We start with the implication from statement (2) to statement (1). Note that $\hat{y}_{G}(p)=1$ implies $\frac{R(G;p)}{S(G;p)}=0$, from which it follows that $R(G;p)=0$.
To see the reverse implication let $G\in\mathcal{H}_{G_{0}}$ be a edge minimal such that $R(G;p)=0$. Now $\hat{y}_{G}(p)=\frac{0}{S(G;p)}+1$. If $S(G;p)\neq 0$, this is well defined and equal to $1$. So let us assume instead $S(G;p)=0$. By minimality of $G$ we may assume by Lemma 2.1 that $G$ is not the series composition of two smaller graphs in $\mathcal{H}_{G_{0}}$. Since by assumption $S(G_{0};p)\neq-R(G_{0};p)$ we have $G\neq G_{0}$ and thus $G$ is the parallel composition of two smaller graphs $G_{1},G_{2}\in\mathcal{H}_{G_{0}}$. Since by Lemma 2.1 $S(G_{1}\parallel G_{2};p)=S(G_{1};p)S(G_{2};p)$, we may assume that $S(G_{1};p)=0$ and moreover we may assume that $G_{2}$ itself is not the parallel composition of two smaller graphs in $\mathcal{H}_{G_{0}}$. By Lemma 2.1 we have
@@
Let $p\in\mathcal{Z}_{G_{0}}$. Then by Lemma 3.1 there exists $G\in\mathcal{H}_{G_{0}}$ such that $\hat{y}_{G}(p)=\frac{R(G;p)}{S(G;p)}+1=1$ and $S(G;p)\neq 0$. We denote $f(p)$ for the map $y\mapsto\hat{y}_{G}(p)$, with $G$ fixed. Since $S(G;p)$ is a polynomial in $p$, we can find $\delta>0$ such that $f|_{B(p,\delta)}$ is holomorphic and so by the open mapping theorem, it is an open map. Let $\varepsilon>0$ such that $\varepsilon<\delta$. Then $f(B(p,\varepsilon))$ is mapped to an open set of $\mathbb{C}$ containing $f(p)=1$, which has to contain some disc $B(1,\rho)$; particularly, it contains the point $1+i\tfrac{\rho}{2}$, which is greater than one in absolute value and not real.

### v2

@@ 1.1. Our contributions
(1.2) ${\mathsf{SP}}(G;p)=\sum_{\begin{subarray}{c}A\subseteq E\\ (V,A)\text{ $s-t$ split}\end{subarray}}(1-p)^{|A|}p^{|E|-|A|}.$
Alternatively, let $\hat{G}$ be the graph obtained from $G$ by merging the vertices $s$ and $t$ (and any edge between $s$ and $t$ becoming a loop). Then ${\mathsf{SP}}(G;p)=R(\hat{G};p)-R(G;p)$. Next we define the effective edge interaction of $G$ at $p$ as
(1.3) $y_{G}(p):=(1-p)\frac{{\mathsf{SP}}(G;p)}{R(G;p)}+1$
@@
and refer to these sets as the zero-locus, density locus and activity locus of $\mathcal{H}_{G_{0}}$ respectively. We moreover define the real analogues of these loci, restricting $p$ to be real, and denote these by $\mathcal{Z}_{G_{0}}^{\mathbb{R}},\mathcal{D}_{G_{0}}^{\mathbb{R}}$ and $\mathcal{A}_{G_{0}}^{\mathbb{R}}$ respectively, where for the real density locus we require the set to be dense in $\mathbb{R}$ and for the real activity locus we also require that $\hat{y}_{G}(p)<-1$ and of course disregard the requirement that $\hat{y}_{G}(p)\notin\mathbb{R}$.
Our first main result relates the different loci defined above and will be proved in Section 3.
@@
$\overline{\mathcal{Z}_{G_{0}}}=\overline{{\mathsf{A}}_{G_{0}}}\quad\text{and}\quad\mathcal{A}_{G_{0}}\setminus\{p\mid R(G_{0};p)=0\}\subseteq\mathcal{D}_{G_{0}}\subseteq\mathcal{A}_{G_{0}}.$
@@
Name: Approx-Abs-Planar-Rel(p)(p)
@@
Name: Approx-Arg-Planar-Rel(p)(p)
@@
We define the problems Approx-Abs-Rel(p)(p) and Approx-Arg-Rel(p)(p) in a similar way, putting no restrictions on the input graph $H$. Note that if $p$ is a real number finding a $0.25$-arg approximation of $R(H;p)$ is equivalent to finding its sign (assuming $R(H;p)\neq 0$).
@@
Let $G_{0}$ be a two-terminal graph. Then for any algebraic number $p\in\mathcal{D}_{G_{0}}\cup\mathcal{D}^{\mathbb{R}}_{G_{0}}$ such that $R(G_{0};p)S(G_{0};p)\neq 0$,
  • the problems Approx-Arg-Rel(p)(p) and Approx-Arg-Rel(p)(p) are #P-hard;
  • if additionally $G_{0}$ is planar with its two terminals on the same face, then the problems Approx-Abs-Planar-Rel(p)(p) and Approx-Arg-Planar-Rel(p)(p) are #P-hard.
@@
For each algebraic number $p\in\mathbb{D}\setminus[0,1)$ both the problems Approx-Abs-Planar-Rel(p)(p) and Approx-Arg-Planar-Rel(p)(p) are #P-hard.
@@ 2.1. Graph operations and reliability
$R(G_{1}\parallel G_{2};p)$ $=R(G_{1};p){\mathsf{SP}}(G_{2};p)\!+\!\!R(G_{2};p){\mathsf{SP}}(G_{1};p)\!+\!\!R(G_{1};p)R(G_{2};p),$
${\mathsf{SP}}(G_{1}\bowtie G_{2};p)$ $=R(G_{1};p){\mathsf{SP}}(G_{2};p)+R(G_{2};p){\mathsf{SP}}(G_{1};p).$
@@
Let $p\in[0,1]$; by definition, this makes $R(G;p)$ and ${\mathsf{SP}}(G;p)$ the probabilities that $G$ remains connected/becomes an sts-t split respectively. We omit the variable $p$ for ease of notation. Then we have the following.
@@ 2.2. Edge interactions
$y_{G_{1}\bowtie G_{2}}$ $=\frac{(1-p){\mathsf{SP}}(G_{1}\bowtie G_{2})}{R(G_{1}\bowtie G_{2})}+1$
$=\frac{(1-p)\left(R(G_{1}){\mathsf{SP}}(G_{2})+R(G_{2}){\mathsf{SP}}(G_{1})\right)}{R(G_{1})R(G_{2})}+1$
$=\frac{(1-p)R(G_{2}){\mathsf{SP}}(G_{1})}{R(G_{1})R(G_{2})}+\frac{(1-p)R(G_{1}){\mathsf{SP}}(G_{2})}{R(G_{1})R(G_{2})}+1$
$=\left(\frac{(1-p){\mathsf{SP}}(G_{1})}{R(G_{1})}+1\right)+\left(\frac{(1-p){\mathsf{SP}}(G_{2})}{R(G_{2})}+1\right)-1$
@@
$\hat{y}_{G_{1}\parallel G_{2}}$ $=\frac{R(G_{1}\parallel G_{2})}{{\mathsf{SP}}(G_{1}\parallel G_{2})}+1$
$=\frac{{\mathsf{SP}}(G_{1}\parallel G_{2})+R(G_{1}\parallel G_{2})}{{\mathsf{SP}}(G_{1}\parallel G_{2})}$
$=\frac{{\mathsf{SP}}(G_{1}){\mathsf{SP}}(G_{2})+R(G_{1}){\mathsf{SP}}(G_{2})+R(G_{2}){\mathsf{SP}}(G_{1})+R(G_{1})R(G_{2})}{{\mathsf{SP}}(G_{1}){\mathsf{SP}}(G_{2})}$
$=\frac{({\mathsf{SP}}(G_{1})+R(G_{1}))({\mathsf{SP}}(G_{2})+R(G_{2}))}{{\mathsf{SP}}(G_{1}){\mathsf{SP}}(G_{2})}$
@@ 2.3. Gadgets
Take a two-terminal graph $(G,s,t)$, which we refer to as a gadget, and another graph $H$. We can substitute the gadget into an edge $e$ of $H$ by removing $e$, adding a copy of $G$, and identifying the source and sink of $G$ with the endpoints of $e$; we denote the resulting graph as $H(G)_{e}$. See Figure 2 below for an example. While technically it matters which endpoint of $e$ is identified with $s$, for the purpose of the reliability polynomial it does not matter as we show below and therefore we don’t specify the choice of identification.
@@
${\mathsf{SP}}(H(G)_{e};p)={\mathsf{SP}}(G;p){\mathsf{SP}}(H\setminus e;p)+R(G;p){\mathsf{SP}}(H/e;p).$
@@
Let $H=(V(H),E(H))$ be a graph and $G=(V(G),E(G))$ a gadget; let $H(G)_{e}=(V,E)$. Consider the structure of a connected edge subgraph of $H(G)_{e}$; every vertex of $G$ has to be connected to at least one of $s,t$ to be connected to the vertices of $H$. A subset $A\subseteq E$ naturally decomposes into $E_{H}\cup E_{G}$, with $A_{H}\subseteq E(H)$, $A_{G}\subseteq E(G)$ and $E_{H}\cap E_{G}=\emptyset$. Then a spanning subgraph $(V,A)$ of $H(G)_{e}$ is connected if and only if one of the following happens:
@@
$\frac{1-p}{R(G;p)}R(H(G)_{e};p)$ $=\frac{1-p}{R(G;p)}\left({\mathsf{SP}}(G;p)R(H\setminus e;p)+R(G;p)R(H/e;p)\right)$
$=(1-p)\frac{{\mathsf{SP}}(G;p)}{R(G;p)}R(H\setminus e;p)+R(H;p)-pR(H\setminus e;p)$
@@
(2.3) $R(H(G);p)$ $=\left(R(G;p)+{\mathsf{SP}}(G;p)\right)^{|E(H)|}R\left(H;\frac{1}{\hat{y}_{G}(p)}\right)\ \ \text{and}$
${\mathsf{SP}}(H(G);p)$ $=\left(R(G;p)+{\mathsf{SP}}(G;p)\right)^{|E(H)|}{\mathsf{SP}}\left(H;\frac{1}{\hat{y}_{G}(p)}\right),$
from which the statement of the lemma directly follows. Observe that $R(G;p)$ is zero as a polynomialproof if and only if $G$ is disconnected. We can then assume that $p\in[0,1]$ is such that $R(G;p)\neq 0$. We work by induction on the number of edges of $H$, starting with the case $|E(H)|=1$. We may assume $H=K_{2}$ (for otherwise the statement is trivial). Since $K_{2}(G)=G$, it is not difficult to see that the desired statement holds. Let us next assume that $|E(H)|>1.$ Let us view $H(G)$ as $H^{\prime}(G)_{e}$ for some edge $e$ of $H$. (Here $H^{\prime}$ is the graph obtained from $H(G)$ by replacing the copy of $G$ on $e$ by $K_{2}$.) By Lemma 2.5 we then have
$R(H(G)$ $;p)=\,R(G;p)R((H/e)(G);p)+{\mathsf{SP}}(G;p)R((H\setminus e)(G);p)$
$=$ $\,\left(R(G;p)+{\mathsf{SP}}(G;p)\right)\!\!\Bigg(\!\!\!\left(1-\frac{1}{\hat{y}_{G}(p)}\right)R((H/e)(G);p)+\frac{1}{\hat{y}_{G}(p)}R((H\setminus e)(G);p)\!\!\Bigg).$
Since $H/e,H\setminus e$ both have fewer edges than $H$, the statement follows by induction and the deletion contraction recurrence. The same steps work if we substitute $R(H(G);p)$ with ${\mathsf{SP}}(H(G);p)$. ∎
@@ 2.4. Multivariate reliability and split reliability
In this section we will define the multivariate analogue of the reliability and split reliability polynomial of a graph. As we will see this will help us to derive a formula for the the general edge replacement, where we replace each edge $e_{i}$ of $H$ with different two-terminal graphs $G_{i}$. For a graph $G=(V,E)$ with edges $\{e_{1},\dots,e_{m}\}$ let $p_{1},\dots,p_{m}$ variables. Then the multivariate reliability polynomial is defined as
@@
Let $H$ be a two-terminal graph with edges $\{e_{1},\dots,e_{m}\}$ and let $G_{1},\dots,G_{m}$ be two-terminal graphs. Let $H(G_{1},\dots,G_{m})$ be a graph obtained from $H$ by replacing, for $i=1,\ldots,m$, the edge $e_{i}$ with the two-terminal graph $G_{i}$ (i.e. we delete the edge $e_{i}=(u,v)$ and we identify the source (resp. target) of $G_{i}$ with $u$ (resp. $v$). Then
@@ 2.5. Normal families and the Montel-Carathèodory Theorem
A family $\mathcal{F}$ of meromorphic functions on an open set $U\subseteq\hat{\mathbb{C}}$ is called normal if every sequence $(f_{n})\subseteq\mathcal{F}$ contains a subsequence which converges spherically uniformly on compact subsets of $U$.
@@
Let $(f_{n})$ be a sequence of meromorphic functions on an open set $U\subseteq\hat{\mathbb{C}}$ which converges spherically uniformly on compact subsets to $f$. Then $f$ is either a meromorphic function on $U$ or identically equal to $\infty$.
@@ 3.1. Relating the activity-locus and the zero-locus
We begin by stating a lemma on the values of the virtual edge interactions (which we recall are defined in (1.4)) at reliability zeros.
@@
Let $p\in\mathbb{C}$ and let $G_{0}$ be a two-terminal graph such that ${\mathsf{SP}}(G_{0};p)\neq-R(G_{0};p)$. Then the following are equivalent:
@@
    $\hat{y}_{G}(p)\in\{\omega,\omega^{2},1\}$, ${\mathsf{SP}}(G;p)\neq 0$ for some $G\in\mathcal{H}_{G_{0}}$ and $\omega$ a primitive third root of unity.
@@
We start with the implication from statement (2) to statement (1). Note that $\hat{y}_{G}(p)=1$ implies $\frac{R(G;p)}{{\mathsf{SP}}(G;p)}=0$, from which it follows that $R(G;p)=0$.
To see the reverse implication let $G\in\mathcal{H}_{G_{0}}$ be a edge minimal such that $R(G;p)=0$. Now $\hat{y}_{G}(p)=\frac{0}{{\mathsf{SP}}(G;p)}+1$. If ${\mathsf{SP}}(G;p)\neq 0$, this is well defined and equal to $1$. So let us assume instead ${\mathsf{SP}}(G;p)=0$. By minimality of $G$ we may assume by Lemma 2.1 that $G$ is not the series composition of two smaller graphs in $\mathcal{H}_{G_{0}}$. Since by assumption ${\mathsf{SP}}(G_{0};p)\neq-R(G_{0};p)$ we have $G\neq G_{0}$ and thus $G$ is the parallel composition of two smaller graphs $G_{1},G_{2}\in\mathcal{H}_{G_{0}}$. Since by Lemma 2.1 $S(G_{1}\parallel G_{2};p)=S(G_{1};p)S(G_{2};p)$, we may assume that $S(G_{1};p)=0$ and moreover we may assume that $G_{2}$ itself is not the parallel composition of two smaller graphs in $\mathcal{H}_{G_{0}}$. By Lemma 2.1 we have
@@
Let $p\in\mathcal{Z}_{G_{0}}$. Then by Lemma 3.1 there exists $G\in\mathcal{H}_{G_{0}}$ such that $\hat{y}_{G}(p)=\frac{R(G;p)}{{\mathsf{SP}}(G;p)}+1=1$ and ${\mathsf{SP}}(G;p)\neq 0$. We denote $f(p)$ for the map $p\mapsto\hat{y}_{G}(p)$, with $G$ fixed. Since ${\mathsf{SP}}(G;p)$ and $R(G;p)$ are distinct polynomials in $p$, we can find $\delta>0$ such that $f|_{B(p,\delta)}$ is holomorphic and not constant and so by the open mapping theorem, it is an open map. Let $\varepsilon>0$ such that $\varepsilon<\delta$. Then $f(B(p,\varepsilon))$ is mapped to an open set of $\mathbb{C}$ containing $f(p)=1$, which has to contain some disc $B(1,\rho)$; particularly, it contains the point $1+i\tfrac{\rho}{2}$, which is greater than one in absolute value and not real.